A man's age is three times his daughters age.Ten years ago,he was five times his daughter's age.Calculate the age of the father.
Answer by stanbon(75887)

A man's age is three times his daughters age.Ten years ago,he was five times his daughter's age.Calculate the age of the father.
-----------------------
Equations:
m = 3d
m-10 = 5(d-10)
----
Substitute for "m" and solve for "d":
3d-10 = 5d-50
2d = 40
d = 20 years (daughter's age)
----
m = 3d = 60 years (man's age)
================
Cheers,
Stan H.
================